uppose that you have a supply of a 20% solution of alcohol and a 80% solution of alcohol. How many quarts of each should be mixed to produce 80 quarts that is 35% alcohol?

Answer by mananth(16949)

Alcohol I 20 ---------------- x quarts
Alcohol II 80 ---------------- 80 - x quarts
Mixture 35.00% ---------------- 80

20x+80(80-x)=35* 80

20x+6400 -80x=2800
20x-80x =2800-6400
-60x= -3600
/ -60
x=60 quarts 20.00% Alcohol I
20 quarts 80.00% Alcohol II